OpenStudy (anonymous):

5 5xtothe3rd __________ 10xtothe7th.....just simplify

OpenStudy (anonymous):

PLEASE HELP

myininaya (myininaya):

\[5x^3/(10x^7)=1/(2x^4)\]

myininaya (myininaya):

5/10=1/2 x^3/x^7=x^(3-7)=x^(-4)=1/(x^4)

OpenStudy (anonymous):

\[7x ^{4} \div5x ^{3}\]